OpenStudy (anonymous):

Two finite sets have m and n elements . The total no. of subsets of the first set is 56 more than the total no.of subsets in the second set. Find the value of m and n .

OpenStudy (anonymous):

let A and B be two sets having m and n numbers of elements respectively number of subsets of A=2m number of subsets of B=2n according to question 2^m-2^n=56 \[2^n (2^{m-n} \ -1)=23(23-1) \] \[\therefore \] \[n=3 \ and \ m-n=3\] \[m=6 \]

OpenStudy (anonymous):

Can you get it now!

OpenStudy (anonymous):

yea .. but can you explain how in 2^n ( 2^m-n - 1) came ? like i know taking 2^n common but why 2^m-n inside ?

OpenStudy (anonymous):

so that it gets cancelled ?

OpenStudy (anonymous):

@IllasMcKay there ?

OpenStudy (anonymous):

ya!

OpenStudy (anonymous):

okay .. ty for the help :)

OpenStudy (anonymous):